# Multi-Agent Deep Reinforcement Learning for Zonal Ancillary Market Coupling

### v1

@@ III-A Nonconvex Game with Side Constraints
III-B Relaxed Generalized Nash Game
In order to extend the results from [17], we consider a relaxed version of our problem, which makes it more tractable and it will allow us to use mild assumptions to obtain an equilibrium existence guarantee. We add the constraints in Eqs. (3) and (6) to the objective function, which gives the function $J_{n}^{p}(w)$, the new constraint maps $h_{n}^{p}(w_{n})$ and $G^{p}(w)$, and the associated set of penalties $M_{n}$ for each agent (see Appendix B). We can formulate the penalized generalized Nash game $\Gamma^{p}$ as follows:
$\min_{w_{n}\in\mathcal{W}_{n}}$ $-J_{n}^{p}(w)$ (15a)
s.t. $h_{n}^{p}(w_{n})\leq 0$ (15b)
$G^{p}(w_{n},w_{-n})\leq 0$ (15c)
Proposition 1.
If $K\geq\frac{Z}{3}$, there exists a set of penalties $(M_{j})_{j\in\mathcal{N}}$ such that a solution of the penalized game (15) is a solution of $\Gamma$.
@@ III-C Existence of a Generalized Nash Equilibrium for $\Gamma$
To obtain existence result for the relaxed problem, we have to verify assumptions on the variables’ domain of definition, the objective functions and the constraint sets.
@@
    There exists $w^{\sharp}\in\mathcal{W}:=\prod_{n\in\mathcal{N}}\mathcal{W}_{n}$ such that $h(w_{n}^{\sharp})<0,\,\forall n$ and $G_{n}(w_{n}^{\sharp},w_{-n}^{\sharp})<0,\,\forall n\in\mathcal{N}$;
  • (d)
    The Hessians matrices $\nabla^{2}h_{n}^{p}(w_{n})$ are copositive;
  • (e)
    The Hessians matrices $\nabla^{2}G_{n}^{p}(w)$ are copositive.
@@
Given $\mathcal{W}$, we assume there exists $\mathcal{W}^{\star}\subseteq\mathcal{W}$, such that for any $w\in\mathcal{W}^{\star}$, the Hessian matrix of the objective function is positive semi-definite. Then, if Assumptions 1 hold, we conclude that inside $\mathcal{W}^{\star}$ the penalized problem has a Nash equilibrium and, consequently, the penalized Generalized Nash game $\Gamma$ also has a Nash equilibrium.
@@
Assumption 1 (a) hold for our settings, since the definition of $\mathcal{W}_{n}$ creates polyhedral and compact sets, i.e., we can express any condition of the form $x\in[x_{a},x_{b}]$ as a set of inequalities $x\leq x_{b},x\geq x_{a}$, which are defined as intersections between closed half spaces.
The objective function is continuous and differentiable everywhere in the domain $\mathcal{W}^{\star}$, and the constraints functions are all bilinear or trilinear. Thus, they are all twice continuous and differentiable over the whole domain $\mathcal{W}^{\star}$. Assumption 1 (b) is then satisfied.
Assumption 1 (c) requires the existence of an interior point. Considering $w^{\sharp}$ that strictly satisfies Eq. (3), (4), (5) and (6), we immediately obtain that $\lambda_{n}^{z}=0,\mu_{n,k}=0,\sigma_{n}=0,\delta_{n}=0,\,\forall k\in[1,K_{n}],\,\forall n\in\mathcal{N},\,\forall z\in\mathcal{Z}$. Eq. (12c) needs instead to be relaxed to a strict inequality.
Finally, for Assumptions 1 (d)-(e), we need to compute the Hessians matrices associated to the constraints functions. Regarding the local set of constraints $h_{n}^{p}(w_{n})$, we obtain $\nabla^{2}h_{n}^{p}(w_{n})=0,\,\forall n\in\mathcal{N},\,\forall w_{n}\in\mathcal{W}_{n}$, except for Eq. (12c), for which the Hessian (we leave the full form in Appendix D) is easily obtained and it satisfies the copositivity requirement. Considering the only remaining global constraint, Eq (5), we obtain a constant positive Hessian (full computation in the Appendix), which directly satisfies the copositivity assumption.
We can now apply [17, Theorem 13] to the $(N+1)-$players game to obtain the existence of a Nash equilibrium. Further, using results from [17, Sec. 2.1], we have an equivalence between the Nash equilibrium of the $(N+1)-$players game and the associated generalized Nash game, which in our case is the penalized version $\Gamma^{p}$. We conclude using Proposition 1, which states the equivalence between $\Gamma^{p}$ and $\Gamma$. ∎
@@
Using [18, Def. 2.1], we have two requirements for a GNG to be a GPG. The first requirement amounts to have a feasible set that is nonempty, which is easily satisfied in our setting. In particular, the feasible set, which we call $C$, does not have to be the cartesian product of the players’ feasible sets, meaning that $C\subseteq\prod_{n=1}^{N}\bar{\mathcal{W}}_{n}(w_{-n})$, where $\bar{\mathcal{W}}_{n}(w_{-n}):=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0,G(w)\leq 0\}$. The second requirement is for a potential $P(w)$ function to exist, which in our case is satisfied by using $P(w):=\sum_{n=1}^{N}J_{n}(w)$. ∎
@@
which can be solved using a standard nonlinear programming solver. The potential structure is also interesting algorithmically, as it can be used to obtain a convergence guarantee for the best-response algorithm.
@@ IV-A Gauss-Seidel Best Response
Furthermore, from [18, Theorem 5.2] and the potential structure of our game, we have a guarantee of convergence to a Nash equilibrium, provided it exists, which we already proved in Section III-C.
@@ IV-B Multi Agent Deep Reinforcement Learning
The intuition behind this algorithm is that the critic learns to map future rewards to specific states and actions (which is the $Q$-function), while the actor tries to maximize the $Q$-function itself (represented by the critic output). What we presented here is the single agent implementation, in our multi-agent settings we use a modified version, where each agent has its own actor and critic, but the critics receive as input the actions of all the agents. This was introduced in [20] as a way of stabilizing the learning and improving the results. In particular, given the zonal nature of our environment, we share the actions between the agents of the same zone, but not between all agents. In Alg. 2 we report in pseudo-code the structure for the training of the RL agents.
We leave further details on the implementation in Appendix E.
@@ V Simulation Results
We now discuss the numerical result of the three algorithmic solutions we propose. We test these algorithms on real data from the German and Austrian ancillary service markets, for the period between January 2024 and August 2024 [21]. The maximum length for the bid sequence is 5, and each submitted capacity must be $\geq 5$ MW. As shown in Tab. II, the demand in Germany is much higher and volatile than that in Austria. Furthermore, the export limit is fixed at $80$ MW for both zones and the core portion is $100$ MW for Austria and 00 MW for Germany. The symmetric and low export limit suggests that producers in Germany could have some impact in Austria, while the converse situation is less likely to arise.
@@
Finally, we test multiple values for the export constraints, in order to study the effect that the two zones have on each other. All simulations were done on a MacBook Pro with an M2 Pro chip and 16G of memory [^2] The code is available at the repository https://github.com/FrancescoMorri/MARL_ancillary_markets.
@@ V-A Convergence of the Multi-Agent RL Simulation
We start by analyzing the convergence and stability of the multi-agent simulation. Our settings are particularly complicated, since the market environment is dependent on all agents decisions. We tested two settings for the MARL algorithm: one with 4 learning based producers, and one with 8. In the first case, the 4 other producers always bid their marginal price. The second case creates a very dynamic environment, since at every step the price of the market is decided based on the actions of 8 independent RL agents. To assess the stability and convergence, we consider the reward of each agent throughout the training. More precisely, we compute the average reward during an episode, this being a full pass through the data. With this data we obtain Fig. 2, where we plot the episodic average reward for each agent divided by zone, considering the case with only RL based producers. Notice that the reward is not the profit, as we mention in the full description of the algorithm in the Appendix. From Fig. 2 we can see that each agent increases its reward during the initial episodes and eventually plateaus to a stable value. These results are particularly interesting, since they show how the algorithm we use is well suited for a medium scale simulation with competitive learning agents, providing results that we can compare with the classical solutions.
@@ V-B MARL Solution and Classical Solutions
Regarding the solution obtained from the MARL algorithm, we have less theoretical guarantees, but we can consider it as a more realistic simulation of the market, since every producer has access only to its own information and its only aim is to increase its profit.
@@ V-C Effects of Coupling Between Zones
We can see that in the case of $4$ RL agents, the average cost is not influenced by the export constraint: this is likely due to the fact that the presence of static agents, always bidding at marginal price, already drives the cost down, and it is not possible to decrease it further. The case with $8$ RL agents is much more interesting: we observe that the cost for Austria is rather stable, while the one for Germany is basically linearly decreasing. We can interpret this behavior considering the differences between the producers in Austria and Germany: in the smaller zone we have less producers, with lower capacities and marginal prices, while in the larger Germany the producers have higher marginal prices. Once the Austrian producers can enter the German market with their leftover capacity, they can bid the same way they would bid for their local market, which in turns, forces the German producers to bid lower, in order to be more competitive. Furthermore, since it is the market that decides how to divide each bid between zone, a producer can not create a priori a specific bid for a specific zone, each bid needs to be general enough to be competitive on every zone available.
We conclude by mentioning [23], where the authors simulated a more economically sophisticated model of the European electricity market, to test what the addition of another smaller zone would do to the costs in the French zone. Their results show the same trend we observe here, with the smaller zone obtaining similar cost, while the larger one decreases its own. Our main methodological advantage, using a MARL based simulation, is that we can quickly test a range of different settings, not just a single scenario, providing a less detailed but broader view of the effect of different market designs. Further, using pre-trained agents, we are testing how a bidding strategy developed for a certain market design translates to a different one, when there is no time to adapt. This approach is of particular interest for situations where the coupling may change due to grid failures, with no warning.
@@ VI Conclusions
In this work we presented a model for multi-zonal ancillary markets based on a bilevel formulation. Each electricity producer is represented by an optimization problem with local and coupling constraints, with the objective of maximizing its profit, while the market operator objective is to minimize the total cost, so we create an adversarial formulation. We then reformulated the bilevel problem as a single level optimization, from the perspective of each producer, using the KKT reformulation of the market optimization. We interpret the ensemble of single level problems of each producer as a generalized Nash game. The game we obtain has side constraints, trilinear objective functions and nonconvex constraints, making it particularly challenging to solve. We analyze this game considering a relaxed version, and we obtain existence results for the Nash equilibrium that we relate to the initial generalized Nash game. We conclude our theoretical analysis showing that the generalized Nash game has a generalized potential structure.
@@ Appendix B Penalized Objective Function
We report here the explicit form for the penalized objective function used in the penalized GNG of Eqs. (15). As described in Section III-B, we relax constraints Eqs. (3)-(6), to obtain:

### v2

@@ III-C Existence of a Generalized Nash Equilibrium for $\Gamma$
To obtain existence result for the penlaized problem, we have to verify assumptions on the variables’ domain, the objective functions and the constraint sets. Thanks to our penalized reformulation, we only need to verify a set of mild assumptions:
@@
    There exists $w^{\sharp}\in\mathcal{W}:=\prod_{n\in\mathcal{N}}\mathcal{W}_{n}$ such that $h_{n}(w_{n}^{\sharp})<0,\,\forall n\in\mathcal{N}$ and $G(w^{\sharp})<0$;
@@
Define $\mathcal{W}^{\star}\subseteq\mathcal{W}$ such that for any $w\in\mathcal{W}^{\star}$, the Hessian matrix of the objective function is positive semi-definite. Then, there exists a generalized Nash equilibrium on $\mathcal{W}^{\star}$ solution of $\Gamma^{p}$.
@@
Assumption 1 (a) holds, since $\mathcal{W}_{n}$ is defined as a polyhedral and compact set, i.e., we can express any condition of the form $x\in[x_{a},x_{b}]$ as a set of inequalities $x\leq x_{b},x\geq x_{a}$, which are defined as intersections between closed half spaces. Further, the objective function is continuous and differentiable everywhere in the domain $\mathcal{W}^{\star}$, and the constraints functions are bilinear or trilinear. Thus, they are twice continuous and differentiable over the whole domain $\mathcal{W}^{\star}$. Assumption 1 (b) is then satisfied. Assumption 1 (c) requires the existence of an interior point. Considering $w^{\sharp}$ that strictly satisfies Eq. (3), (4), (5) and (6), we immediately obtain that $\lambda_{n}^{z}=0,\mu_{n,k}=0,\sigma_{n}=0,\delta_{n}=0,\,\forall k\in[1,K_{n}],\,\forall n\in\mathcal{N},\,\forall z\in\mathcal{Z}$. Eq. (12c) needs instead to be relaxed to a strict inequality. We then apply [18, Theorem 13] to the $(N+1)-$players game to obtain the existence of a generalized Nash equilibrium for $\Gamma^{p}$. ∎
Corollary 1.
The generalized Nash game $\Gamma$ defined in (13) admits a solution.
Proof.
Using results from [18, Sec. 2.1], we have an equivalence between the Nash equilibrium of the $(N+1)-$players game and the associated generalized Nash game, which in our case is the penalized version $\Gamma^{p}$. We conclude using Proposition 1, which states the equivalence between $\Gamma^{p}$ and $\Gamma$. ∎
@@
Using [19, Def. 2.1], two requirements need to be satisfied for a GNG to be a GPG. The first requirement amounts to have a feasible set that is nonempty, which is easily satisfied in our setting. In particular, the feasible set, which we call $C$, does not have to be the cartesian product of the players’ feasible sets, meaning that $C\subseteq\prod_{n=1}^{N}\bar{\mathcal{W}}_{n}(w_{-n})$, where $\bar{\mathcal{W}}_{n}(w_{-n}):=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0,G(w)\leq 0\}$. The second requirement is for a potential $P(w)$ function to exist, which in our case is satisfied by using $P(w):=\sum_{n=1}^{N}J_{n}(w)$. ∎
@@
which can be solved using a standard nonlinear programming solver.
@@ IV-A Gauss-Seidel Best Response
Furthermore, from [19, Theorem 5.2] and the potential structure of our game, we have a guarantee of convergence to a Nash equilibrium, provided it exists, which we already proved in Section III-C.
@@ IV-B Multi Agent Deep Reinforcement Learning
The intuition behind this algorithm is that the critic learns to map future rewards to specific states and actions (which is the $Q$-function), while the actor tries to maximize the $Q$-function itself (represented by the critic output). In the multi-agent settings we appply a modified version of the single agent implementation presented here, where each agent has its own actor and critic, but the critics receive as input the actions of all the agents. This was introduced in [21] as a way of stabilizing the learning and improving the results. In particular, given the zonal nature of our environment, we share the actions between the agents of the same zone, but not between all agents. In Alg. 2 we report in pseudo-code the structure for the training of the RL agents.
We leave further details on the implementation and references to the code in Appendix E.
@@ V Simulation Results
We now discuss the numerical results of the three algorithmic solutions we propose. We test these algorithms on real data from the German and Austrian ancillary service markets, for the period between January 2024 and August 2024 [22]. The maximum length for the bid sequence is 5, and each submitted capacity must be $\geq 5$ MW. As shown in Tab. II, the demand in Germany is much higher and volatile than that in Austria. Furthermore, the export limit is fixed at $80$ MW for both zones and the core portion is $100$ MW for Austria and 00 MW for Germany. The symmetric and low export limit suggests that producers in Germany could have some impact in Austria, while the converse situation is less likely to arise.
@@
Finally, we test multiple values for the export constraints, in order to study the effect that the two zones have on each other. All simulations were done on a MacBook Pro with an M2 Pro chip and 16G of memory.
@@ V-A Convergence of the Multi-Agent RL Simulation
We start by analyzing the convergence and stability of the multi-agent simulation. Our settings are particularly complicated, since the market environment is dependent on all agents decisions. We tested two settings for the MARL algorithm: one with 4 learning based producers, and one with 8. In the first case, the 4 other producers always bid their marginal price. The second case creates a very dynamic environment, since at every step the price of the market is decided based on the actions of 8 independent RL agents. To assess the stability and convergence, we consider the reward of each agent throughout the training. More precisely, we compute the average reward during an episode, this being a full pass through the data. With this data we obtain Fig. 2, where we plot the episodic average reward for each agent divided by zone, considering the case with only RL based producers. Notice that the reward is not the profit, as we mention in the full description of the algorithm in the Appendix. From Fig. 2 we can see that each agent increases its reward during the initial episodes and eventually plateaus to a stable value. These results are particularly interesting, since they show how the algorithm is well suited for a medium scale simulation with competitive learning agents, providing results that we can compare with the classical solutions.
We want to stress here that, using the MARL algorithm, we compute a stationary solution for the market, without a guarantee of it being an equilibrium.
@@ V-B MARL Solution and Classical Solutions
Regarding the solution obtained from the MARL algorithm, even if there are less theoretical guarantees, we can consider it as a more realistic simulation of the market, since every producer has access only to its own information and its only aim is to increase its profit.
@@ V-C Effects of Coupling Between Zones
We can see that in the case of $4$ RL agents, the average cost is not influenced by the export constraint: this is likely due to the fact that the presence of static agents, always bidding at marginal price, already drives the cost down, and it is not possible to decrease it further. The case with $8$ RL agents is much more interesting: we observe that the cost for Austria is rather stable, while the one for Germany is basically linearly decreasing. We can interpret this behavior considering the differences between the producers in Austria and Germany: in the smaller zone we have less producers, with lower capacities and marginal prices, while in the larger Germany the producers have higher marginal prices. Once the Austrian producers can enter the German market with their leftover capacity, they can bid the same way they would bid for their local market, which in turns, forces the German producers to bid lower, in order to be more competitive. Furthermore, since it is the market that decides how to divide each bid between zone, a producer can not create a priori a specific bid for a specific zone, each bid needs to be general enough to be competitive on every zone available. We conclude by mentioning [24], where the authors simulated a more economically sophisticated model of the European electricity market, to test what the addition of another smaller zone would do to the costs in the French zone. Their results show the same trend we observe here, with the smaller zone obtaining similar cost, while the larger one decreases its own. Our main methodological advantage, using a MARL based simulation, is that we can quickly test a range of different settings, not just a single scenario, providing a less detailed but broader view of the effect of different market designs. Further, using pre-trained agents, we are testing how a bidding strategy developed for a certain market design translates to a different one, when there is no time to adapt. This approach is of particular interest for situations where the coupling may change due to grid failures, with no warning.
@@ VI Conclusions
In this work we presented a model for multi-zonal ancillary markets based on a bilevel formulation. Each electricity producer is represented by an optimization problem with local and coupling constraints, with the objective of maximizing its profit, while the market operator objective is to minimize the total cost, so we create an adversarial formulation. We then reformulated the bilevel problem as a single level optimization, from the perspective of each producer, using the KKT reformulation of the market optimization. We interpret the ensemble of single level problems of each producer as a generalized Nash game. The game we obtain has side constraints, trilinear objective functions and nonconvex constraints, making it particularly challenging to solve. We analyze this game considering a penlaized version, and we obtain existence results for the Nash equilibrium that we relate to the initial generalized Nash game. We conclude our theoretical analysis showing that the generalized Nash game has a generalized potential structure.
@@ Appendix B Penalized Objective Function
We report here the explicit form for the penalized objective function used in the penalized GNG of Eqs. (14). As described in Section III-A, we relax constraints Eqs. (3)-(6), to obtain: